OpenStudy (anonymous):

Domain and range for these types of graphs? Pictures below. Medal will be awarded!

OpenStudy (anonymous):

How do I find the domain and range here?

OpenStudy (mertsj):

Look at the graph and see what x values and y values are on the graph. x values = domain, y values = range.

OpenStudy (jack1):

domain = all the possible x values range = all the possible y values

OpenStudy (anonymous):

Domain is simply asking, where do the x values start and end, well the x's start at -2 and they seem to go off to the right (infitinity) so the domain is \[x \ge-2 \] or ;2, inf)

OpenStudy (anonymous):

[2, inf) * and the range would be done in the same procudure and the next question is the same thing

OpenStudy (anonymous):

I got the range is \[y \ge-2\] as well but I do not know how to find the range on the next question at all. It is an absolute value then it stops at different numbers? I am confused.

OpenStudy (anonymous):

For the second problem's domain I got all real numbers though.

OpenStudy (anonymous):

The range is never more than 6 though? So would the range be written as \[6\le y \ge0\] ??

OpenStudy (jack1):

yep, sorry, that's right, my bad

OpenStudy (anonymous):

Oh okay! Thanks for your help mate!

OpenStudy (jack1):

nah dude, it was all you that time hey ;)